OpenStudy (lolo_rose):

Find v*w if v=-3i+4j-5k and w=3i-2j+3k

OpenStudy (mrcoolguy):

post in math

OpenStudy (astrophysics):

It's a dot product, so \[v=-3i+4j-5k \] and \[w = 3i-2j+3k\] \[v \cdot w = (-3 \times 3)i +(4 \times -2 ) j + (-5 \times 3)k\]

OpenStudy (astrophysics):

@lolo_rose

OpenStudy (lolo_rose):

thanks!!!